OpenStudy (anonymous):

find the acute angel formed by the two lines. y=1/3x+7 and y=-1/2x-8 please help..

OpenStudy (e.mccormick):

Ah ha. Yah. You can find the angles each of these makes with the x axis. Then, you use this information to know what angles they will have between them. One pair will be the obtuse, the other the acute.

OpenStudy (anonymous):

im confused..?

OpenStudy (e.mccormick):

OK. Well, slope is rise over run, right? Well, what is tangent?

OpenStudy (anonymous):

tan is rise over run?

OpenStudy (e.mccormick):

Yes, it is sine over cosine, so Y over X, which is Rise over Run. That means you can use Tan to find both the angles of each line with respect to the X axis. Then it is math to get them with repect to each other.

OpenStudy (anonymous):

do i plug that into my calculator?

OpenStudy (e.mccormick):

Well, they probably want exact values, so you may need to find out what it is in terms of radian and pi.

OpenStudy (e.mccormick):

I am going to look and see if someone has an equation to solve it faster. But here is what I mean so far in a graphical way.

OpenStudy (e.mccormick):

Take a look at : http://www.transtutors.com/math-homework-help/straight-line/angle-between-two-lines.aspx and see if that makes sense. If not, what part so we can go from there.

OpenStudy (anonymous):

i just tried and te acute angel they make is 98 degrees. i think i did it wrong..?

OpenStudy (anonymous):

when i do tan(1/2) i get 55 degrees and tan(1/3) is 35 degrres when i add them together i get 90 degrees??

OpenStudy (e.mccormick):

Well, that is not the best description. I found another, that is clearer. It is the same thing, but better written. I tested it, and confirmed grapically. http://www.tpub.com/math2/5.htm I'll put in some of the math in another post so you can see what I did.

OpenStudy (e.mccormick):

The key is the: \[\tan \theta = \frac{ m_2-m_1 }{1+ m_2m_1}\] Lets take your slops as m1=-1/2, and m2=1/3.

OpenStudy (e.mccormick):

Then: \[\tan \theta = \frac{ \frac{1}{3}-(-\frac{1}{2}) }{1+ \frac{1}{3}(-\frac{1}{2}) }\]

OpenStudy (e.mccormick):

Now, can you see how that simplifies? It becomes a reaaaaaaaaly easy number.

OpenStudy (anonymous):

0/1? or just 0? or 1?

OpenStudy (anonymous):

is it 1?

OpenStudy (anonymous):

do i do inverse tan(1) ?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

Yah, absolutly! You are there.

OpenStudy (e.mccormick):

The arctan(1)

OpenStudy (anonymous):

so the answer is 78 degrees?

OpenStudy (e.mccormick):

Your calculator is in radians mode.

OpenStudy (anonymous):

no sorry 45 degrees?

OpenStudy (e.mccormick):

Yep! Once I found a better copy of the formula, it was easy. Sorry about linking the poorly done one. It was jsut confusing.

OpenStudy (anonymous):

its okay thank you so so much! you wanna help me with another problem? im gonna post it.

OpenStudy (e.mccormick):

More trig? Sure. I need some trig brushup.

OpenStudy (anonymous):

yep;)

OpenStudy (e.mccormick):

Yah, best way to remember these thing is to work them through and talk about them. So it helps me as much as it helps you.